A woman of mass 41 kg jumps off the bow of
a 59 kg canoe that is initially at rest.
If her velocity is 4.5 m/s to the right, what
is the velocity of the canoe after she jumps?
Answer in units of m/sˆ
Don't round answer.
1
Expert's answer
2019-12-04T08:38:59-0500

The law of conservation of momentum gives


"p_i=p_f"

"0=mv-Mu"

The velocity of the canoe


"u=\frac{mv}{M}=\frac{41\times 4.5}{59}=3.127\:\rm m/s"